What is the probability of choosing a queen a king or an ace from a standard deck of playing cards?
The probability of drawing a queen, king, or ace from a standard deck of 52 cards is 12 in 52, or 3 in 13, or about 0.2308.
What is the probability of not getting a 7 on the roll of two dice?
Find the probability of getting a sum of 7 and call that p(7). Now the complement rule says 1-p(7) is the probability of not getting 7.
The ways we get 7 are 1 and 6, or 2 and 5, or 3 and 4.
If you make a table with 1-6 on the top and 1-6 on the side, you can let each entry in your table be the sum of the top row and the side column. Since this is a 6x6 table, you can see
there are 36 outcomes.
6 of those outcome result in a sum of 7.
They are:
1+6
6+1
2+5
5+2
3+4
4+3
So p(7) =6/36 or 1/6. P of not getting 7 is 1-1/6=5/6
What is the probability of rolling a six-sided dice twice and coming up with at least one number 5?
First calculate the probability of NOT getting a single 5. This probability is 5/6 x 5/6 = 25/36. Therefore, the probability of getting at least one 5 is the complement thereof, that is, 1 - 25/36 = 11/36.
How do you write a unit rate in word form?
it depends on the unit rate you are using but here are some examples
mph= miles per hour
kph= kilometres per hour
fps= feet per second
cps= centimetres per second
What is the probability of selecting a spade face card from a deck of 52?
3 in 52 (jack, queen, king of spades)
The answer depends on what the underlying probability distribution function is.
Most statistical textbooks will contain tables for the binomial, Poisson, Standard Normal, Student's t, Fisher's F, Chi-squared, Pearson's r tests. Then there are lots of non-parametric tests (Spearman's Rank, Mann Whitney U, Kolmogorov-Smirnoff for example) for which there are other tables but these are less widely published. And lastly, there are tests for even more esoteric distributions.
It is normal for the ear canal to have some yellowish-brown earwax
The events in an experiment are mutually exclusive if only one can occur at a time true or false?
provide one business-related example each, with explanation, for mutually exclusive and independent events
What is the probability of rolling 3 3s?
Assuming a standard 6-sided die with numbers 1-6, the probability is:
(1/6) * (1/6) * (1/6) = 1/216 = 0.004629 = 0.4629%
If you are rolling more than three dice, the odds improve dramatically.
What constitutes probable cause to search a vehicle in Utah?
A reasonable belief that a crime has been committed. Example: You pull over a vehicle and while speaking with the driver smell the odor of marijuana in the vehicle.
Added: Or observe a firearm, or what appears to be proceeds of crime inside the vehicle.
ALSO: Be advised that these examples are only a small sample of reasonable causes to suspect criminal activity.
What is the probability of getting a black face card from a well shuffled pack of card?
There are 6 black face cards in a standard deck of 52 cards, so the probability of drawing a black face card in one try is 6/52 = 0.115
What is the importance of unexpected outcomes?
depends on what subject: for psychology, stress is elevated more when people are suprised with new unpredicted challenges rather than handling a known stress.
In an experiment, it would mean the variables are not related and another confounding varialbe is affecting the change, statistical indifference?
12%
What do the four kings in card deck represent?
French Playing Cards Got Their Face Card Designs because they were named after, and designed to look like, actual historical figures.
The King of Hearts was Charlemagne;
the King of Diamonds was Julius Caesar;
the King of Clubs was Alexander the Great,
and the King of Spades was King David from the Holy Bible.
Do identical twins with opposite genders exist?
Yes, it's possible for a mutation in the chromosome to produce otherwise-identical twins with opposite genders. But generally no, identical twins have the same gender. (but fraternal twins have different genders just as often as any siblings)
What is the probability that jesus is the messiah?
It is said in Isaiah that he will be the Messiah, and in another chapter that he would die for the sins of the world.
Roughly: 1/2 * 1/2 * 1/2 = 1/8.
To get an exact answer, we'd need to know whether these guesses are completely random, or if the guesser is using a strategy (e.g. if the first was black, is the guesser going to choose red because there are more red than black cards left or not). Using a strategy would make the guesser get them all right slightly more than 1/8th of the time (the exact odds can be calculated). Guessing randomly will, on average, make you right 1/8th of the time.
Do you eat more when your having a boy then a girl?
There is no evidence to support the hypothesis that you will eat more when you are having a boy then a girl as opposed to having a girl and then a boy.
300,000,000,000 m/s
How many students must be in a room the be sure at least two are of the same sex?
Assuming no student is intersex, the answer is 3.
What is lowest hand in 5 card stud?
The lowest hand in standard poker is 2, 3, 4, 5, and 7, not in the same suit.
What does the word rank mean in data sets?
It means the position [of an observation] in an ordered data set.
A dime is toss three time probability that it will come up head all three times is .125?
Yes, if it is a fair coin.
What are the uses of probability in computer science?
Good cos that's what I bin doin!On one site the board with the numbers on is disalpyed back to front , i.e. with the number 1 on the right hand side.So I still place my chips starting the sequence from the left hand end, and moving them as per the worksheet, so although the actual numbers differ, the position of the chips stay the same as shown in the worksheet.If you get what I mean! Same thing but different .